- Press Release
- Dec 9, 2022
NASA Space Station On-Orbit Status 18 February 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
After wake-up, morning hygiene and breakfast, the crew worked in the Progress-351 (16P) vehicle, dismantling the Kurs-A rendezvous and approach radar system of its SUDN motion control & navigation system and removing it from the transport drone, a three-hour job. These valuable components will be returned to Earth for reuse. [KURS-A is the active half of the Russian space program’s proven S-band radar system for automated flight, which measures relative motion parameters between Progress (or Soyuz) and the ISS during rendezvous operations, to enable the autopilot’s calculation of corrective impulses. The system’s passive transponder counterpart (KURS-P) is on the Service Module (SM), with one antenna each at the tip of the two solar array wings.]
Afterwards, FE Sharipov worked ~1.5 hours in the DC-1 “Pirs” docking module, removing and replacing one of two auxiliary digital commutator (DKTs TA253B) electronics boxes, a subsystem of the ASS message acquisition system. The old device was discarded.
Using new uplinked procedural instructions, CDR/SO Chiao investigated three failed GLAs (General Luminaire Assemblies), one in the Node, two in the Lab. [By identifying failure signatures, verifying FDI LED (Fault Detection Isolation light-emitting diode) status, photographing any physical markings or damage and repairing the dimmer switch if required, the procedure is intended to assist the ground in identification of the root cause of the relatively frequent GLA failures.]
On the Russian SOZh/ECLSS environment control and life support system, Sharipov removed and replaced the gas/liquid separator unit in the no. 2 condensate separation & pumping subsystem (BRPK-2) with a new unit. [The removed unit was stowed in the SM.]
Leroy Chiao retrieved the spectrometer and detector components of the TEPC (Tissue Equivalent Proportional Counter) from their temporary location in the Node and relocated them to the SM’s starboard crew quarters area, with the detector actually inside the crew quarters on the deck. [On 1/31, Leroy had moved the TEPC to the Node from its original position in the Lab. Since the spectrometer is now again plugged in a regular CHeCS power/data outlet, this time on an SM panel, the ground can again receive automated TEPC telemetry.]
Salizhan performed the periodic (weekly) replenishing of the Elektron system’s water supply for electrolysis, his 14th, filling the KOV thermal loops EDV container with water from SM Rodnik tank #2 (BV2). [The procedure is specially designed to prevent air bubbles from getting into the BZh liquid unit where their pressure spikes, from collapsing, could cause micropump impeller cavitation and Elektron shutdown, as happened numerous times in the past. Usually, when the Rodnik’s potable water is not used, the EDV water is condensate drawn from the BKO multifiltration/purification column and the air/liquid separator unit (GZhS) while the crewmember checks for any air bubbles in the EDV (and, if visible, estimates their number). Elektron water is also supplied from USOS condensate in a CWC (collapsible water container) that is checked for its contents of air bubbles and is rejected if the estimated total air bubble volume is more than 30 cubic centimeters (1 cm air bubble is about 0.5 ccm).]
At ~6:25am EST, Sharipov conducted his regular weekly IMS tagup with specialists at TsUP/Moscow, discussing open issues concerning identification of equipment and storage locations for the IMS databases. [Today’s topics included storage location for a PK-3 (Plasma Crystal) kit, listings of Soyuz air duct components, and new equipment relocations this week.]
Salizhan attended to the routine daily maintenance of the SOZh environment control & life support system, while Leroy prepared the regular IMS (Inventory Management System) “delta” file for the daily automated export/import to the three IMS databases on the ground. Chiao also performed a procedure on the onboard IMS database to restore it after some corruption that happened on 1/13.
The CDR transferred the hard-disk drive (HDD) of the medical equipment computer (MEC) into the chassis of the old SSC-1 computer that now takes over MEC duties. [The SSC-1 chassis (#6075), a spare after its replacement with an A31p laptop on 1/31, could not be used as a PCS (portable computer system) machine, unlike the former MEC chassis (#6077), which now becomes a PCS.]
On request from the ground, Chiao did some additional loads testing on UOP-4 (Utility Outlet Panel 4) that had been under investigation after the trip of RPC-4 (Remote Power Controller #4) in a Lab RPCM (RPC Module) on 11/11/04.
During its search in the electronic box (EB) of the TVIS treadmill yesterday, the missing battery did not turn up. The treadmill was returned to its nominal configuration in the SM floor “pit”, and a 10-minute speed characterization test was performed by Leroy today. [Absence of the EB battery has to date not affected TVIS operations negatively. Its purpose is to “remember” time and date as well as particular relay positions if TVIS power is cycled off/on at the SM circuit breakers or further upstream, which is not generally the case. If it should happen, resetting of time/date and control panel LEDs would be done manually.]
The crew worked out in accordance with their daily 2.5-hr. physical exercise protocol on TVIS (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (aerobic plus anaerobic). [Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 1 of a new set) and one hour on VELO.]
Leroy then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
Working off the voluntary Russian task list, Salizhan performed the regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment, which researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.
At ~2:40pm, Chiao and Sharipov are scheduled for their regular (nominally weekly) teleconference with the ISS Flight Director at MCC-Houston via S-band/audio.
Overnight, Progress 16 propellants from the cargo ship’s refueling system (SD) and remaining propellants from its KDU fuel and oxidizer tanks were transferred to ISS storage in the FGB module. [Monitored by Moscow on Daily Orbits 3-5, the SM’s automated daily timeline sequencer (SPP) commanded transition to propellant transfer mode, then commenced propellant transfer from the prop tanks of Progress to the four tanks of the FGB via prop lines passing through the SM, lasting about 1-1.5 hrs.]
Battery reconditioning on BCDU (battery charge/discharge unit) 4B3 is proceeding as planned. [The battery set will complete its second of two complete discharge cycles on 2/20 (Sunday). Once completely discharged, the batteries will be recharged to 100% capacity and brought back online.]
Yesterday’s on-board fire emergency drill went smoothly, with good coordination of the crew with MCC-H and TsUP/Moscow.
The crew was congratulated by PAO on their excellent TV downlink for the ESA Space Conference in Brussels, Belgium, yesterday. Their words on the importance and strength of the international partnership were both timely and well received by the assembled dignitaries.
Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Mt. Kilimanjaro, Tanzania (Dynamic Event. This near-nadir pass over Mt. Kilimanjaro presented an opportunity for photography of the summit glaciers. The high elevation of the summit [5895 m] frequently keeps it exposed above surrounding clouds. Frequent imagery of the summit glaciers is important as they are receding at a high rate and are predicted to vanish over the next twenty years. Looking to the right of track for the summit; it was most likely surrounded by clouds), and Patagonian Glaciers, S. America (this nadir pass over the northern Patagonian ice fields provided an opportunity for detailed mapping. Overlapping image swaths across the smaller mountain glaciers are useful for monitoring of changes in ice volume and extent. The smaller glaciers tend to be more sensitive indicators of climate change due to quicker response to cooling or warming trends).
CEO photography can be viewed and studied at the websites:
See also the website “Space Station Challenge” at:
To view the latest photos taken by the expedition 10 crew visit:
- http://spaceflight.nasa.gov/gallery/images/station/crew-10/ndxpage1.html at NASA’s Human Spaceflight website.
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
- Progress M-51 (16P) undocking & destructive reentry — 2/26/05;
- Progress M-52 (17P) launch — 2/28/05.
- EVA-13 — 3/25/05;
- Soyuz TMA-6 (10S) launch — 4/15/05 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
- Soyuz TMA-5 (9S) undock — 4/25/05 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
- Progress M-53 (18P) launch — 6/10/05;
- Progress M-54 (19P) launch — 8/24/05;
- Soyuz TMA-7 (11S) launch — 9/27/05.
ISS Orbit (as of this morning, 7:18am EST [= epoch]):
- Mean altitude — 358.9 km
- Apogee height — 362.0 km
- Perigee height — 355.9 km
- Period — 91.72 min.
- Inclination (to Equator) — 51.64 deg
- Eccentricity — 0.000456
- Solar Beta Angle — -4.8 deg (magnitude decreasing, until tomorrow)
- rbits per 24-hr. day — 15.70
- Mean altitude loss in last 24 hours — 180 m
- Revolutions since FGB/Zarya launch (Nov. 98) — 35702
ISS Altitude History
Apogee height — Mean Altitude — Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.